# Fighting Exponentially Small Gaps by Counterdiabatic Driving

### v2

@@ Appendix F Derivation of the minimization equations for the Max Cut
Appendix G Derivation of the variational equations with COLD for the Max Cut
In this section, we provide a detailed derivation of how the COLD approach can be involved in the local variational expansion for the Max Cut problem. We apply a longitudinal field as the optimizable operator with time-dependent coefficient given by the Fourier expansion as
$f=\sum_{k=1}^{N_{k}}\gamma_{k}\sin(k\pi t/T),$ (192)
with $N_{k}$ denoting the number of considered Fourier modes. In this case, the full Hamiltonian first and second-order CD corrections read as
$H^{(2)}_{\mathrm{COLD}}=-\lambda\sum_{c}\frac{1-\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2}-(1-\lambda)\sum_{j}\sigma^{x}_{j}+\sum_{j}\epsilon_{j}\sigma^{y}_{j}+\sum_{c}\alpha_{c}(\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}})+f\sum_{j}\sigma^{z}_{j}.$ (193)
The integral action reads
$G(H^{(2)}_{\mathrm{COLD}})=-\sum_{c=1}^{M}\frac{1+8(1-\lambda)\alpha_{c}}{2}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}+\sum_{c=1}^{M}\frac{1+2\lambda\alpha_{c}}{2}(\sigma^{x}_{c_{1}}+\sigma^{x}_{c_{2}})+4(1-\lambda)\sum_{c=1}^{M}\alpha_{c}\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}$
$+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left[(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{1},c^{\prime}_{2}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\right]$
$-f\sum_{c}\alpha_{c}(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}+\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}})-f\sum_{j}\epsilon_{j}\sigma^{x}_{j}+\partial_{\lambda}f\sum_{j}\sigma^{z}_{j}-(1-\lambda)\sum_{j}\epsilon_{j}\sigma^{z}_{j}-\frac{\lambda}{2}\sum_{c}(\epsilon_{c_{2}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}+\epsilon_{c_{1}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}})$
$=-\sum_{c=1}^{M}\frac{1+8(1-\lambda)\alpha_{c}}{2}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}+\sum_{c=1}^{M}\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)\sigma^{x}_{c_{1}}+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)\sigma^{x}_{c_{2}}+4(1-\lambda)\sum_{c=1}^{M}\alpha_{c}\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}$
$+\partial_{\lambda}f\sum_{j}\sigma^{z}_{j}+\lambda\sum_{c,c^{\prime}}\alpha_{c}\left[(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}+\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{1},c^{\prime}_{2}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\right]$
$-\sum_{c}\left((f\alpha_{c}+\lambda\epsilon_{c_{1}}/2)\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}+(f\alpha_{c}+\lambda\epsilon_{c_{2}}/2)\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\right)+\sum_{j}\left[\partial_{\lambda}f-(1-\lambda)\epsilon_{j}\right]\sigma^{z}_{j}.$ (194)
By similar techniques to those in the previous section, the action to be minimized becomes
$+\sum_{c,c^{\prime}}\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+f\frac{\epsilon_{c^{\prime}_{1}}}{3}\right)\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+f\frac{\epsilon_{c^{\prime}_{2}}}{3}\right)(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}$
$+\sum_{c,c^{\prime}}\left[\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)^{2}+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)^{2}\right]\delta_{c_{1},c^{\prime}_{1}}\delta_{c_{2},c^{\prime}_{2}}$
$+\sum_{c,c^{\prime}}\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+f\frac{\epsilon_{c^{\prime}_{2}}}{3}\right)\delta_{c_{1},c^{\prime}_{2}}+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)\left(\frac{1+2\lambda\alpha_{c^{\prime}}}{2}+f\frac{\epsilon_{c^{\prime}_{1}}}{3}\right)\delta_{c_{2},c^{\prime}_{1}}$
$+\lambda^{2}\sum_{c,c^{\prime}}(\alpha_{c}+\alpha_{c^{\prime}})^{2}\delta_{c_{1},c^{\prime}_{2}}+\lambda^{2}\sum_{c,c^{\prime}}\alpha^{2}_{c}\left[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})+(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{c_{2},c^{\prime}_{2}}\right]+16(1-\lambda)^{2}\sum_{c}\alpha^{2}_{c}+\sum_{j}\left[\partial_{\lambda}f-(1-\lambda)\epsilon_{j}\right]^{2}$
$+\sum_{c}(f\alpha_{c}+\lambda\epsilon_{c_{1}}/2)^{2}+(f\alpha_{c}+\lambda\epsilon_{c_{2}}/2)^{2}.$ (195)
The corresponding minimization equations are obtained by taking the variations with respect to both $\epsilon_{j}$ and $\alpha_{c}$
$\frac{\delta\mathrm{Tr}\left[G^{\dagger}\,G\right]}{\delta\epsilon_{n}}=$ (196)
$\frac{2}{3}f\sum_{c,c(n)}\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)\delta_{n,c_{1}}(1-\delta_{c_{2}(n),c_{2}})+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)(1-\delta_{c_{1}(n),c_{1}})\delta_{n,c_{2}}$
$+\sum_{c(n)}\frac{4}{3}f\left(\frac{1+2\lambda\alpha_{c(n)}}{2}+f\frac{\epsilon_{n}}{3}\right)+\frac{2}{3}f\sum_{c,c(n)}\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{2}}}{3}\right)\delta_{n,c_{2}}+\left(\frac{1+2\lambda\alpha_{c}}{2}+f\frac{\epsilon_{c_{1}}}{3}\right)\delta_{n,c_{1}}$
$+\lambda\sum_{c(n)}(f\alpha_{c(n)}+\lambda\epsilon_{n}/2)-2(1-\lambda)(\partial_{\lambda}f-(1-\lambda)\epsilon_{n})=0,$
where $c(n)$ denotes the edge in which the $n$-th spin appears and the summation over holds for every edge in which $n=c_{1}$ or $n=c_{2}$. Rearranging it so that all $\epsilon_{n}$-s appear on the l.h.s., one obtains
$\frac{2}{9}f^{2}\sum_{c,c(n)}\epsilon_{c_{1}}(\delta_{n,c_{2}}+\delta_{n,c_{1}}(1-\delta_{c_{2}(n),c_{2}}))+\epsilon_{c_{2}}(\delta_{n,c_{1}}+(1-\delta_{c_{1}(n),c_{1}})\delta_{n,c_{2}})+\frac{2}{3}\left[2f^{2}+\lambda^{2}+6(1-\lambda)^{2}\right]\epsilon_{n}$ (197)
$+\frac{2}{3}f\lambda\sum_{c,c(n)}\alpha_{c}\left[(1-\delta_{c_{1}(n),c_{1}})\delta_{n,c_{2}}+\delta_{n,c_{1}}(1-\delta_{c_{2}(n),c_{2}})+\delta_{n,c_{1}}+\delta_{n,c_{2}}\right]+\frac{7}{3}\lambda f\sum_{c(n)}\alpha_{c(n)}$
$=2\partial_{\lambda}f(1-\lambda)-(\frac{4}{3}fL+4+2fL).$
The minimization with respect to $\alpha_{c}$ reads as
[…]

### v3

@@ Appendix F Derivation of the minimization equations for the Max Cut
Appendix G Gap amplification within the local CD expansion
In this section, we provide further details about the gap amplification of the local CD expansion in the Max Cut and $3$-XORSAT models. The relative difference energy gaps in the bare process without CD and in that assisted by local CD admits a remarkably universal form of $(\Delta^{(1,2)}_{\mathrm{min,CD}}-\Delta_{\mathrm{min}})/\Delta_{\mathrm{min}}\sim T^{-2}$ for both models, as shown in Fig. 18 and in Fig. 19.
Figure 18: Relative difference between ground-state gaps without CD and with the local CD approaches. The improvement achieved by first-order (solid lines, filled dots) and second-order (dashed lines, empty dots) local expansions exhibits an approximate quadratic decay with system size. Notably, the amplification of the gap vanishes at driving times just preceding the onset of finite fidelity.
Figure 19: Relative difference between the energy gaps without CD and with the first-order ansatz decaying approximately as the minimal spin glass bottleneck independently of the system size. No essential gap amplification is present around the driving times, where even without CD, the fidelity departs from zero.
Appendix H Derivation of the variational equations with COLD for the Max Cut
In this section, we derive the equations for the optimized COLD approach with the second-order commutator as optimizable operators. Due to the structure of these extra operators, the general form of the local CD up to three-body terms takes the same form as that of the second-order local expansion. First, we write out explicitly the optimizable operator,
$\mathcal{O}_{\mathrm{opt}}$ $=$ $\beta^{(1)}\left[H_{0},[H_{1},H_{0}]\right]+\beta^{(2)}\left[H_{1},[H_{1},H_{0}]\right]$ (180)
$=$ $\beta^{(1)}\,\sin(2\pi\lambda)\,\Bigg[3\sum_{j=1}^{L}\sigma^{x}_{j}+\sum_{c,c^{\prime}}\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\,\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}$
$\quad+\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{1},c^{\prime}_{1}})\,\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}+2\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\,\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\Bigg]$
$+2\beta^{(2)}\,\sin(2\pi\lambda)\,\sum_{c}\Big(\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}-\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}\Big),$
where $H_{1}=\sum_{j=1}^{L}\sigma^{x}_{j}$ and $H_{P}=\sum_{c=1}^{M}\frac{1+\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}}{2}$ is the problem Hamiltonian, Eq. (30), for the Max Cut, respectively. From this, the full Hamiltonian reads
$H^{(2)}_{\mathrm{COLD}}$ $=$ $\frac{\lambda+4\beta^{(2)}\sin(2\pi\lambda)}{2}\sum_{c}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}-\big((1-\lambda)-3\beta^{(1)}\sin(2\pi\lambda)\big)\sum_{j}\sigma^{x}_{j}-2\beta^{(2)}\sin(2\pi\lambda)\sum_{c}\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}$ (181)
$+2\beta^{(1)}\sin(2\pi\lambda)\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\,\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}$
$+\beta^{(1)}\sin(2\pi\lambda)\sum_{c,c^{\prime}}\Big[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\,\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{1},c^{\prime}_{1}})\,\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}+2\delta_{c_{2},c^{\prime}_{1}}\,\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\Big]$
$+\sum_{c}\alpha_{c}\big(\sigma^{z}_{c_{1}}\sigma^{y}_{c_{2}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}\big)+\sum_{c,c^{\prime}}\beta_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\big(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{c^{\prime}_{2}}+\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\big).$
The corresponding $G$ operator measuring the deviations from the exact CD,
$G=H_{1}-H_{P}+\partial_{\lambda}\mathcal{O}_{\mathrm{opt}}-i\left[-\lambda H_{p}-(1-\lambda)H_{1}+\mathcal{O}_{\mathrm{opt}},H^{(1)}_{1}+H^{(2)}_{1}\right$], (182)
can be written in the full explicit form as
$G$ $=$ $\sum_{j=1}^{L}\sigma^{x}_{j}-\frac{1}{2}\sum_{c}1+\left(-\frac{1}{2}+4\pi\cos(2\pi\lambda)\beta^{(2)}\right)\sum_{c}\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}-4\pi\cos(2\pi\lambda)\beta^{(2)}\sum_{c}\sigma^{y}_{c_{1}}\sigma^{y}_{c_{2}}$
$+2\pi\cos(2\pi\lambda)\,\beta^{(1)}\Bigg[3\sum_{j=1}^{L}\sigma^{x}_{j}+\sum_{c,c^{\prime}}\delta_{c_{1},c^{\prime}_{1}}\big(1-\delta_{c_{2},c^{\prime}_{2}}\big)\,\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}+\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{2}}\big(1-\delta_{c_{1},c^{\prime}_{1}}\big)\,\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c_{2}}+2\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\,\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\Bigg]$
$-\lambda\sum_{c,c^{\prime}}\alpha_{c}\,\delta_{c_{2},c^{\prime}_{1}}\,\big(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{c^{\prime}_{2}}+\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\big)-\lambda\sum_{c,c^{\prime}}\alpha_{c}\,\delta_{c_{1},c^{\prime}_{1}}\big(1-\delta_{c_{2},c^{\prime}_{2}}\big)\,\big(\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{y}_{c^{\prime}_{2}}+\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{x}_{c^{\prime}_{2}}\big)$
$-\lambda\sum_{c,c^{\prime}}\alpha_{c}\,\delta_{c_{2},c^{\prime}_{2}}\big(1-\delta_{c_{1},c^{\prime}_{1}}\big)\,\big(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{c^{\prime}_{1}}+\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{1}}\big)-2\lambda\sum_{c,c^{\prime},c^{\prime\prime}}\beta_{c,c^{\prime}}\,\delta_{c^{\prime}_{2},c^{\prime\prime}_{1}}\,\big(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{c^{\prime\prime}_{2}}+\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime\prime}_{2}}\big)$
$-2(1-\lambda)\sum_{c,c^{\prime}}\beta_{c,c^{\prime}}\,\delta_{c_{2},c^{\prime}_{1}}\,\big(\sigma^{x}_{c_{2}}\sigma^{y}_{c_{1}}\sigma^{z}_{c^{\prime}_{2}}+\sigma^{x}_{c_{1}}\sigma^{y}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\big)$
$-2\,\beta^{(1)}\,\sin(2\pi\lambda)\Bigg[\sum_{j=1}^{L}\sum_{c}\alpha_{c}\Big(\delta_{j,c_{1}}(\sigma^{y}_{j}\sigma^{y}_{c_{2}}-\sigma^{z}_{j}\sigma^{z}_{c_{2}})+\delta_{j,c_{2}}(\sigma^{y}_{j}\sigma^{y}_{c_{1}}-\sigma^{z}_{j}\sigma^{z}_{c_{1}})\Big)$
$+\sum_{c,c^{\prime}}\sum_{d}\alpha_{d}\Big(\delta_{c_{1},d_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\,\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{y}_{d_{2}}-\delta_{c_{1},d_{2}}(1-\delta_{c_{2},c^{\prime}_{2}})\,\sigma^{y}_{d_{1}}\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}$
$+\delta_{c_{2},d_{2}}(1-\delta_{c_{1},c^{\prime}_{1}})\,\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{y}_{c_{2}}\sigma^{y}_{d_{1}}-\delta_{c_{2},d_{1}}(1-\delta_{c_{1},c^{\prime}_{1}})\,\sigma^{y}_{d_{2}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{y}_{c_{2}}$
$+2\delta_{c_{2},c^{\prime}_{1}}\delta_{c_{1},d_{1}}\,\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{y}_{d_{2}}-2\delta_{c_{2},c^{\prime}_{1}}\delta_{c_{1},d_{2}}\,\sigma^{y}_{d_{1}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\Big)$
$+\sum_{c,c^{\prime}}\sum_{d,d^{\prime}}\beta_{d,d^{\prime}}\,\delta_{d_{2},d^{\prime}_{1}}\Big(\delta_{c_{1},d_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\,\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{x}_{d_{2}}\sigma^{y}_{d^{\prime}_{2}}-\delta_{c_{1},d_{2}}(1-\delta_{c_{2},c^{\prime}_{2}})\,\sigma^{y}_{d_{1}}\sigma^{y}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{x}_{d_{2}}\sigma^{y}_{d^{\prime}_{2}}$
$+\delta_{c_{2},d_{2}}(1-\delta_{c_{1},c^{\prime}_{1}})\,\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{y}_{c_{2}}\sigma^{x}_{d_{1}}\sigma^{y}_{d^{\prime}_{2}}-\delta_{c_{2},d_{1}}(1-\delta_{c_{1},c^{\prime}_{1}})\,\sigma^{y}_{d_{2}}\sigma^{z}_{c_{1}}\sigma^{z}_{c^{\prime}_{1}}\sigma^{y}_{c_{2}}\sigma^{x}_{d_{1}}\sigma^{y}_{d^{\prime}_{2}}$
$+2\delta_{c_{2},c^{\prime}_{1}}\delta_{c_{1},d_{1}}\,\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{x}_{d_{2}}\sigma^{y}_{d^{\prime}_{2}}-2\delta_{c_{2},c^{\prime}_{1}}\delta_{c_{1},d_{2}}\,\sigma^{y}_{d_{1}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{x}_{d_{2}}\sigma^{y}_{d^{\prime}_{2}}\Big)\Bigg]$
$-4\,\beta^{(2)}\,\sin(2\pi\lambda)\Bigg[\sum_{c}\sum_{d}\alpha_{d}\Big(\delta_{c_{1},d_{1}}(\sigma^{x}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{y}_{d_{2}}-\sigma^{x}_{c_{1}}\sigma^{y}_{c_{2}}\sigma^{z}_{d_{2}})+\delta_{c_{2},d_{2}}(\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{y}_{d_{1}}-\sigma^{y}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{d_{1}})\Big)$
$+\sum_{c,c^{\prime}}\sum_{d,d^{\prime}}\beta_{d,d^{\prime}}\,\delta_{d_{2},d^{\prime}_{1}}\Big(\delta_{c_{1},d_{1}}\,\sigma^{x}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{y}_{d^{\prime}_{2}}-\delta_{c_{1},d_{2}}\,\sigma^{y}_{d_{1}}\sigma^{x}_{c_{2}}\sigma^{x}_{c_{1}}\sigma^{z}_{c^{\prime}_{2}}\sigma^{z}_{d^{\prime}_{2}}$
$+\delta_{c_{2},d_{2}}\,\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{x}_{c^{\prime}_{1}}\sigma^{y}_{d_{1}}\sigma^{y}_{d^{\prime}_{2}}-\delta_{c_{2},d_{1}}\,\sigma^{y}_{d_{2}}\sigma^{z}_{c_{1}}\sigma^{x}_{c_{2}}\sigma^{x}_{c^{\prime}_{1}}\sigma^{y}_{d^{\prime}_{2}}\Big)\Bigg].$
The corresponding action is given by the same terms as in the second-order case, supplemented by the proper additional terms. Denoting again the terms from the second-order expansion by $(1-\lambda)^{2}S_{X}$ and $\lambda^{2}S_{ZZ}$, one obtains
$+16(1-\lambda)^{2}\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\beta^{2}_{c,c^{\prime}}+\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left[4(1-\lambda)\beta_{c,c^{\prime}}-\lambda(\alpha_{c}+\alpha_{c^{\prime}})\right]^{2}+\lambda^{2}\sum_{c,c^{\prime},c^{\prime\prime}}\left(\delta_{c^{\prime}_{2},c_{1}}\beta^{2}_{c^{\prime},c}\delta_{c^{\prime}_{1},c^{\prime\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\beta^{2}_{c,c^{\prime}}\delta_{c^{\prime\prime}_{1},c^{\prime}_{2}}\right)$
$+\sum_{c}\left(\left[-\tfrac{1}{2}+4\pi\cos(2\pi\lambda)\beta^{(2)}\right]^{2}+\left[-4\pi\cos(2\pi\lambda)\beta^{(2)}\right]^{2}\right)$
$+\big(6\beta^{(1)}\sin(2\pi\lambda)\big)^{2}\sum_{c,d}\alpha_{c}\alpha_{d}\big(\delta_{c_{1},d_{1}}+\delta_{c_{1},d_{2}}+\delta_{c_{2},d_{1}}+\delta_{c_{2},d_{2}}\big)$ (185)
$+\big(\beta^{(1)}\sin(2\pi\lambda)\big)^{2}\sum_{c,c^{\prime},d,d^{\prime}}\alpha_{c^{\prime}}\alpha_{d^{\prime}}\Big[\delta_{c_{1},c^{\prime}_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})\delta_{d_{1},d^{\prime}_{1}}(1-\delta_{d_{2},d^{\prime}_{2}})(\delta_{c_{1},d_{1}}\delta_{c_{2},d_{2}}\delta_{c^{\prime}_{2},d^{\prime}_{2}}+\delta_{c_{2},d_{1}}\delta_{c_{1},d_{2}}\delta_{c^{\prime}_{2},d^{\prime}_{2}}+\delta_{c^{\prime}_{2},d_{1}}\delta_{c_{2},d_{2}}\delta_{c_{1},d^{\prime}_{2}})$
$+\delta_{c_{2},c^{\prime}_{2}}(1-\delta_{c_{1},c^{\prime}_{1}})\delta_{d_{2},d^{\prime}_{2}}(1-\delta_{d_{1},d^{\prime}_{1}})(\delta_{c_{1},d_{1}}\delta_{c^{\prime}_{1},d^{\prime}_{1}}\delta_{c_{2},d_{2}}+\delta_{c_{1},d_{2}}\delta_{c^{\prime}_{1},d^{\prime}_{1}}\delta_{c_{2},d_{1}}+\delta_{c^{\prime}_{1},d_{2}}\delta_{c_{1},d_{1}}\delta_{c_{2},d^{\prime}_{1}})\Big]$
$+\big(\beta^{(1)}\sin(2\pi\lambda)\big)^{2}\sum_{c,c^{\prime},d,d^{\prime}}\beta_{c,c^{\prime}}\beta_{d,d^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\delta_{d_{2},d^{\prime}_{1}}\big(\delta_{c_{1},d_{1}}\delta_{c_{2},d_{2}}\delta_{c^{\prime}_{2},d^{\prime}_{2}}+\delta_{c_{1},d_{2}}\delta_{c_{2},d_{1}}\delta_{c^{\prime}_{2},d^{\prime}_{2}}\big)$
$+\big(2\beta^{(2)}\sin(2\pi\lambda)\big)^{2}\sum_{c,d}\alpha_{c}\alpha_{d}\big(\delta_{c_{1},d_{1}}\delta_{c_{2},d_{2}}+\delta_{c_{1},d_{2}}\delta_{c_{2},d_{1}}\big)$
$+\big(2\beta^{(2)}\sin(2\pi\lambda)\big)^{2}\sum_{c,c^{\prime},d,d^{\prime}}\beta_{c,c^{\prime}}\beta_{d,d^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\delta_{d_{2},d^{\prime}_{1}}\big(\delta_{c_{1},d_{1}}\delta_{c_{2},d_{2}}+\delta_{c_{1},d_{2}}\delta_{c_{2},d_{1}}\big)$
$+2\big(\beta^{(1)}\sin(2\pi\lambda)\big)\big(2\beta^{(2)}\sin(2\pi\lambda)\big)\sum_{c,c^{\prime},d,d^{\prime}}\beta_{c,c^{\prime}}\beta_{d,d^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\delta_{d_{2},d^{\prime}_{1}}\big(\delta_{c_{1},d_{1}}\delta_{c_{2},d_{2}}+\delta_{c_{1},d_{2}}\delta_{c_{2},d_{1}}\big)$
$+\sum_{c,c^{\prime},d,d^{\prime}}\beta_{c,c^{\prime}}\beta_{d,d^{\prime}}\,2\,\beta^{(1)}\sin(2\pi\lambda)\,\delta_{c_{2},c^{\prime}_{1}}\delta_{d_{2},d^{\prime}_{1}}\Big[\delta_{c_{1},d_{1}}(1-\delta_{c_{2},c^{\prime}_{2}})(1-\delta_{d_{2},d^{\prime}_{2}})+\delta_{c_{2},d_{2}}(1-\delta_{c_{1},c^{\prime}_{1}})(1-\delta_{d_{1},d^{\prime}_{1}})$
$+2\delta_{c_{2},c^{\prime}_{1}}\delta_{c_{1},d_{1}}\delta_{d_{1},d^{\prime}_{2}}-2\delta_{c_{2},c^{\prime}_{1}}\delta_{c_{1},d_{2}}\delta_{d_{1},d^{\prime}_{2}}\Big]$
$+\sum_{c,d,d^{\prime}}\alpha_{c}\beta_{d,d^{\prime}}\,2\,\beta^{(1)}\sin(2\pi\lambda)\Big[\delta_{c_{1},d_{1}}(1-\delta_{c_{2},d^{\prime}_{2}})+\delta_{c_{2},d_{2}}(1-\delta_{c_{1},d^{\prime}_{1}})+\delta_{c_{1},d_{2}}+\delta_{c_{2},d_{1}}\Big]+\sum_{c}\alpha_{c}\,6\,\beta^{(1)}\sin(2\pi\lambda)$
$+\sum_{c,c^{\prime}}\beta_{c,c^{\prime}}\,2\,\beta^{(1)}\sin(2\pi\lambda)\,\delta_{c_{2},c^{\prime}_{1}}+\sum_{c,d}\alpha_{c}\alpha_{d}\,2\,\beta^{(2)}\sin(2\pi\lambda)\Big[\delta_{c_{1},d_{1}}+\delta_{c_{2},d_{2}}-\delta_{c_{1},d_{2}}-\delta_{c_{2},d_{1}}\Big]$
$+\sum_{c,c^{\prime},d,d^{\prime}}\beta_{c,c^{\prime}}\beta_{d,d^{\prime}}\,2\,\beta^{(2)}\sin(2\pi\lambda)\,\delta_{c_{2},c^{\prime}_{1}}\delta_{d_{2},d^{\prime}_{1}}\Big[\delta_{c_{1},d_{1}}+\delta_{c_{2},d_{2}}-\delta_{c_{1},d_{2}}-\delta_{c_{2},d_{1}}\Big]$
[…]